### v3

@@ 4.2 Description of the algorithm
drawn from distributions $P$ and $Q$, each being a mixture of $m$ multivariate Gaussian or $t$-distributions. Let $P_{\ell}$ and $Q_{r}$ be the empirical distributions of $\aleph_{1}$ and $\aleph_{2}$. Applying Theorems 5 and 6, the discrepancy between the clusters is measured by the average of the projected KS-distances between $P_{\ell}$ and $Q_{r}$ along a strong sm-uniqueness set of random directions (Theorem 7).
@@
1:Integers $m,d$; samples $\aleph_{1}=\{X_{1},\dots,X_{\ell}\}\subset\mathbb{R}^{d}$, $\aleph_{2}=\{Y_{1},\dots,Y_{r}\}\subset\mathbb{R}^{d}$; fix $k\geq\tfrac{1}{2}(2m-1)(d^{2}+d-2)-1$
@@
Theorems 5 and 6 show that if this limit is zero, then $P=Q$. However, the asymptotic distribution of $\mathrm{MA}_{k}$ is unknown, since projections onto different directions are only conditionally independent.
@@ 4.3 Simulations
Fig. 5: Simulation per Scenario is provided considering $\eta_{1}=1$, $\eta_{2}=2$, and $\eta_{3}=1/2$ in Scenarios 1, 2, and 3, respectively
@@
Fig. 6: Boxplot of $500$ replicates of $\textit{MA}_{k}$ (Right panels) and $D_{k}$ (Left panels), with $k=100$, calculated between mixtures $F_{1}$ and $F_{2}$ by varying $\eta_{1},\eta_{2}$ and $\eta_{3}$ in the scenarios.
In Figure 5, for improved visualization, one representative simulation per scenario is displayed, corresponding to $\eta_{1}=1$, $\eta_{2}=2$, and $\eta_{3}=1/2$ in Scenarios 1, 2, and 3, respectively. The maximum aggregation measure $D_{k}$ quantifies the largest observed separation between the mixtures $F_{1}$ and $F_{2}$ across the $k$ nearest-neighbor comparisons. This measure is particularly sensitive to pronounced discrepancies (whether in location, dispersion, or component weights), thereby providing a robust indicator of distinct distributional features.
@@ Appendix A Algorithm 1: Methodological Aspects
This appendix provides a detailed account of the steps in Algorithm 1.
A.1 Step 1: Estimation of mixture weights
For each random direction $u$, the projected distribution is
$\sum_{j=1}^{n}\lambda_{j}N\!\Bigl{(}\langle u,\mu_{j}\rangle,\langle u,\Sigma_{j}u\rangle\Bigr{)},$
with characteristic function
$\psi_{u}(t)=\sum_{j=1}^{n}\lambda_{j}\exp\!\left(i\langle u,\mu_{j}\rangle t-\tfrac{1}{2}\langle u,\Sigma_{j}u\rangle t^{2}\right).$
Its empirical counterpart is
$\widehat{\psi}_{u}(t)=\frac{1}{N}\sum_{j=1}^{N}\exp(i\langle u,X_{j}\rangle t).$
Let $t_{1},\dots,t_{m}$ be fixed grid points. Define
$Z_{u}$ $:=\bigl{(}\operatorname{Re}\psi_{u}(t_{1}),\dots,\operatorname{Re}\psi_{u}(t_{m}),\operatorname{Im}\psi_{u}(t_{1}),\dots,\operatorname{Im}\psi_{u}(t_{m})\bigr{)}^{T},$
$F_{u}(\upsilon,\sigma^{2},\Lambda)$ $:=\bigl{(}\operatorname{Re}\widehat{\psi}_{u}(t_{1}),\dots,\operatorname{Re}\widehat{\psi}_{u}(t_{m}),\operatorname{Im}\widehat{\psi}_{u}(t_{1}),\dots,\operatorname{Im}\widehat{\psi}_{u}(t_{m})\bigr{)}^{T}.$
The estimator $(\widehat{\upsilon}_{u,(1)},\widehat{\sigma}^{2}_{u,(1)},\widehat{\Lambda}_{u,(1)})$ minimizes
$\bigl{(}Z_{u}-F_{u}(\upsilon,\sigma^{2},\Lambda)\bigr{)}^{T}W\bigl{(}Z_{u}-F_{u}(\upsilon,\sigma^{2},\Lambda)\bigr{)},$
where $W$ is a weighting matrix, estimated for instance using HAC methods [58]. This corresponds to a generalized method-of-moments estimator [42]. The procedure is replicated in $k$ directions $u_{1},\dots,u_{k}$, and
$\widehat{\Lambda}_{(1)}=\frac{1}{k}\sum_{r=1}^{k}\widehat{\Lambda}_{u_{r},(1)}.$
A.2 Step 2: Estimation of means and covariances
Fixing $\Lambda=\widehat{\Lambda}_{(1)}$, repeat the estimation in each $u_{r}$ for $(\upsilon,\sigma^{2})$. Denote the resulting estimators as $\widehat{\upsilon}_{u_{r},(2)}=(\widehat{\upsilon}_{u_{r},(2),1},\dots,\widehat{\upsilon}_{u_{r},(2),n})$ and $\widehat{\sigma}^{2}_{u_{r},(2)}=(\widehat{\sigma}^{2}_{u_{r},(2),1},\dots,\widehat{\sigma}^{2}_{u_{r},(2),n})$.
Then reconstruct the original parameters by solving
$\widehat{\mu}_{j}$ $=\operatorname*{argmin}_{\mu_{j}}\sum_{r=1}^{k}\Bigl{(}\langle u_{r},\mu_{j}\rangle-\widehat{\upsilon}_{u_{r},(2),j}\Bigr{)}^{2},$ (6)
$\widehat{\Sigma}_{j}$ $=\operatorname*{argmin}_{\Sigma_{j}\in S^{+,d}}\sum_{r=1}^{k}\Bigl{(}\langle u_{r},\Sigma_{j}u_{r}\rangle-\widehat{\sigma}^{2}_{u_{r},(2),j}\Bigr{)}^{2},$ (7)
for $j=1,\dots,n$, with $S^{+,d}$ the cone of $d\times d$ positive semidefinite matrices.
Acknowledgments
Fraiman and Moreno were supported by grant FCE-3-2022-1-172289 from ANII (Uruguay), Program for the Development of Basic Sciences-PEDEClBA (Uruguay) and 22520220100031UD from CSIC (Uruguay). Ransford was supported by NSERC Discovery Grant RGPIN-2020-04263.

### v4

@@ 3.5 An example with real data
An important aspect of this example is the comparison of student performance in acquired mathematical skills between those attending private schools and those attending public schools. The comparison considers two variables: the score assigned to the item and the socioeconomic and cultural level of the child’s home. Comparing the means $\widehat{\mu}_{1}=(-0.14,-0.15)$ for those attending public schools with $\widehat{\mu}_{2}=(1.16,0.47)$ for those attending private schools, we see a clear drop in the performance of students attending public schools, particularly marked in the first coordinate, but also significant in the second. This highlights a problem to be considered by the country’s policymakers.
@@ 4.1 Introduction
Our proposal does not, in principle, provide distances between partitions directly. Instead, it provides a measure between two model-based distributions associated with the partitions. More precisely, in our framework, the objects being compared are the mixture distributions associated with the two partitions (i.e., the distributions induced by the model-based clusterings), instead of the label assignments themselves. For this reason, we report $D_{k}$ and $\mathrm{MA}_{k}$ defined below as descriptive measures of agreement between partitions.
@@ 4.2 Description of the algorithm
drawn from distributions $P$ and $Q$, each being a mixture of $m$ multivariate Gaussian or $t$-distributions. Let $P_{\ell}$ and $Q_{r}$ be the empirical distributions of $\aleph_{1}$ and $\aleph_{2}$. Applying Theorems 4 and 5, the discrepancy between the clusters is measured by the average of the projected KS-distances between $P_{\ell}$ and $Q_{r}$ along a strong sm-uniqueness set of random directions (Theorem 6).
@@
1:Integers $m,d$; samples $\aleph_{1}=\{X_{1},\dots,X_{\ell}\}\subset\mathbb{R}^{d}$, $\aleph_{2}=\{Y_{1},\dots,Y_{r}\}\subset\mathbb{R}^{d}$; fix $k\geq\tfrac{1}{2}(2m-1)(d^{2}+d-2)+1$
@@
Theorems 4 and 5 show that if this limit is zero, then $P=Q$. However, the asymptotic distribution of $\mathrm{MA}_{k}$ is unknown, since projections onto different directions are only conditionally independent.
@@
(iii) Pre-whitening via $\Sigma^{-1/2}$ is a convenient option to mitigate scale (and unit-of-measure) effects when the practitioner wishes to control for them; in that case, one may apply a whitening transform to the data before computing the projected KS aggregations.
@@ 4.3 Simulations
Fig. 8: Representative realization in each scenario comparing mixtures $F_{1}$ and $F_{2}$: Scenario 1 varies the location shift $\eta_{1}$ (with $\eta_{2}=\eta_{3}=0$), Scenario 2 varies the scale factor $\eta_{2}$ (with $\eta_{1}=\eta_{3}=0$), and Scenario 3 varies the mixing-weight perturbation $\eta_{3}$ (with $\eta_{1}=\eta_{2}=0$).
@@
Fig. 9: Agreement between mixtures $F_{1}$ and $F_{2}$ using random projections: boxplots of $D_{k}=\max_{1\leq i\leq k}\mathrm{KS}(i)$ (left) and $\mathrm{MA}_{k}=\frac{1}{k}\sum_{i=1}^{k}\mathrm{KS}(i)$ (right), over 500 replicates with $k=100$ directions and 500 draws from each mixture per replicate. Scenario 1 varies location ($\eta_{1}$), Scenario 2 varies scale ($\eta_{2}$), and Scenario 3 varies mixing weights ($\eta_{3}$).
In Figure 8, for improved visualization, one representative simulation per scenario is displayed, corresponding to $\eta_{1}=1$, $\eta_{2}=2$, and $\eta_{3}=1/2$ in Scenarios 1, 2, and 3, respectively. The maximum aggregation measure $D_{k}$ quantifies the largest observed separation between the mixtures $F_{1}$ and $F_{2}$ across the $k$ projected KS comparisons. This measure is particularly sensitive to pronounced discrepancies (whether in location, dispersion, or component weights), thereby providing a robust indicator of distinct distributional features.
@@ 5 Conclusions and future work
The two parts of the manuscript are linked by a common projection-based representation for Gaussian/$t$-mixtures: the same finite family of directions that supports identifiability and reconstruction in Section 3 also yields a natural distributional notion of partition agreement in Section 4, obtained by aggregating one-dimensional discrepancies across directions.
@@
Acknowledgments
We are grateful to the four anonymous referees for their thoughtful suggestions and detailed feedback, which led to substantial improvements in the presentation and overall quality of the manuscript.
Fraiman and Moreno were supported by grant FCE-3-2022-1-172289 from ANII (Uruguay), the Program for the Development of Basic Sciences–PEDECIBA (Uruguay), and 22520220100031UD from CSIC (Uruguay). Ransford was supported by NSERC Discovery Grant RGPIN-2020-04263.
@@ Appendix A Algorithm 1: Methodological aspects
This appendix details the implementation of Algorithm 1.
A.1 Step 1: Estimation of the mixing weights via one-dimensional ECF fits
Let $P$ be an $m$-component Gaussian mixture on $\mathbb{R}^{d}$ with parameters $\Theta=(\Lambda,\bm{\mu},\bm{\Sigma})$, where $\Lambda=(\lambda_{1},\dots,\lambda_{m})$, $\sum_{j=1}^{m}\lambda_{j}=1$, $\bm{\mu}=(\mu_{1},\dots,\mu_{m})$, and $\bm{\Sigma}=(\Sigma_{1},\dots,\Sigma_{m})$.
Fix a unit direction $u\in\mathbb{S}^{d-1}$ and define the projected data
$Y_{i}(u):=\langle u,X_{i}\rangle,\qquad i=1,\dots,N.$
Under the model, $Y_{i}(u)$ follows a univariate Gaussian mixture with component-specific projected means and variances
$\upsilon_{u,j}:=\langle u,\mu_{j}\rangle,\qquad\sigma^{2}_{u,j}:=\langle u,\Sigma_{j}u\rangle,\qquad j=1,\dots,m,$
that is,
$P_{\langle u\rangle}=\sum_{j=1}^{m}\lambda_{j}\,\mathcal{N}\!\big(\upsilon_{u,j},\sigma^{2}_{u,j}\big).$
Characteristic functions.
The (model) characteristic function of the projected mixture is
$\psi_{u}(t;\upsilon_{u},\sigma_{u}^{2},\Lambda):=\sum_{j=1}^{m}\lambda_{j}\exp\!\left(it\,\upsilon_{u,j}-\tfrac{1}{2}t^{2}\,\sigma^{2}_{u,j}\right),$
where $\upsilon_{u}=(\upsilon_{u,1},\dots,\upsilon_{u,m})$ and $\sigma_{u}^{2}=(\sigma^{2}_{u,1},\dots,\sigma^{2}_{u,m})$. Its empirical counterpart (ECF) is
$\widehat{\psi}_{u,N}(t):=\frac{1}{N}\sum_{i=1}^{N}\exp\!\big(it\,Y_{i}(u)\big)=\frac{1}{N}\sum_{i=1}^{N}\exp\!\big(it\,\langle u,X_{i}\rangle\big).$
Grid and moment vector.
Fix grid points $t_{1},\dots,t_{M}$ (e.g., $t_{\ell}=\tau\,\ell$ as in [32]). Define the empirical moment vector
$\widehat{Z}_{u}:=\big(\operatorname{Re}\widehat{\psi}_{u,N}(t_{1}),\dots,\operatorname{Re}\widehat{\psi}_{u,N}(t_{M}),\ \operatorname{Im}\widehat{\psi}_{u,N}(t_{1}),\dots,\operatorname{Im}\widehat{\psi}_{u,N}(t_{M})\big)^{\top},$
and the model moment vector
$Z_{u}$ $(\upsilon_{u},\sigma_{u}^{2},\Lambda)$
$:=\big(\operatorname{Re}\psi_{u}(t_{1};\upsilon_{u},\sigma_{u}^{2},\Lambda),\dots,\operatorname{Re}\psi_{u}(t_{M};\cdot),\ \operatorname{Im}\psi_{u}(t_{1};\cdot),\dots,\operatorname{Im}\psi_{u}(t_{M};\cdot)\big)^{\top}.$
Weighted least-squares / GMM criterion.
For a positive semidefinite weighting matrix $W\in\mathbb{R}^{2M\times 2M}$, we estimate $(\upsilon_{u},\sigma_{u}^{2},\Lambda)$ by minimizing
$Q_{u}(\upsilon_{u},\sigma_{u}^{2},\Lambda):=\big(\widehat{Z}_{u}-Z_{u}(\upsilon_{u},\sigma_{u}^{2},\Lambda)\big)^{\top}W\big(\widehat{Z}_{u}-Z_{u}(\upsilon_{u},\sigma_{u}^{2},\Lambda)\big),$ (6)
over $\Lambda$ in the simplex and $\sigma^{2}_{u,j}>0$. Following [76, 82], $W$ may be estimated using HAC-type procedures (e.g. [58]), yielding a generalized method-of-moments estimator in the sense of [42].
Multiple directions and label alignment.
Draw $u_{1},\dots,u_{k}\sim\mathrm{Unif}(\mathbb{S}^{d-1})$ independently and compute, for each $r=1,\dots,k$, the one-dimensional estimate
$(\widehat{\upsilon}_{u_{r},(1)},\widehat{\sigma}^{2}_{u_{r},(1)},\widehat{\Lambda}_{u_{r},(1)}).$
Since each projected fit is performed independently, component labels may be permuted across directions. We align labels as follows.
Choose a pivot direction $u_{r^{\ast}}$ (e.g. the one maximizing the separation between the estimated projected means) and order its components by increasing $\widehat{\upsilon}_{u_{r^{\ast}},(1),j}$. For each $r$, match the $m$ triples $(\widehat{\lambda}_{u_{r},(1),j},\widehat{\upsilon}_{u_{r},(1),j},\widehat{\sigma}^{2}_{u_{r},(1),j})$ to the pivot triples by solving a minimum-cost assignment problem (e.g. Hungarian algorithm), producing aligned estimates $(\widetilde{\upsilon}_{u_{r},(1)},\widetilde{\sigma}^{2}_{u_{r},(1)},\widetilde{\Lambda}_{u_{r},(1)})$. Finally, define the Step 1 weight estimator by averaging:
$\widehat{\Lambda}_{(1)}:=\frac{1}{k}\sum_{r=1}^{k}\widetilde{\Lambda}_{u_{r},(1)}.$ (7)
Remark (Student $t$-mixtures).
For $t$-mixtures, the projected model remains a univariate $t$-mixture with projected locations $\upsilon_{u,j}=\langle u,\mu_{j}\rangle$ and projected scales determined by $\langle u,\Sigma_{j}u\rangle$; Step 1 is identical, replacing $\psi_{u}(\cdot)$ by the corresponding $t$-mixture characteristic function.
A.2 Step 2: Re-estimation of projected means/variances and reconstruction in $\mathbb{R}^{d}$
With $\Lambda$ fixed at $\widehat{\Lambda}_{(1)}$, we re-estimate, for each direction $u_{r}$, the projected location and variance parameters by minimizing the same criterion (6) over $(\upsilon_{u},\sigma_{u}^{2})$ only. Denote the resulting aligned estimators by
$\widehat{\upsilon}_{u_{r},(2)}=(\widehat{\upsilon}_{u_{r},(2),1},\dots,\widehat{\upsilon}_{u_{r},(2),m}),\qquad\widehat{\sigma}^{2}_{u_{r},(2)}=(\widehat{\sigma}^{2}_{u_{r},(2),1},\dots,\widehat{\sigma}^{2}_{u_{r},(2),m}).$
Reconstruction of mean vectors.
For each component $j=1,\dots,m$, we reconstruct $\mu_{j}\in\mathbb{R}^{d}$ via least squares:
$\widehat{\mu}_{j}:=\operatorname*{argmin}_{\mu\in\mathbb{R}^{d}}\ \sum_{r=1}^{k}\big(\langle u_{r},\mu\rangle-\widehat{\upsilon}_{u_{r},(2),j}\big)^{2}.$ (8)
Let $U\in\mathbb{R}^{k\times d}$ be the matrix with $r$-th row $u_{r}^{\top}$ and let $\widehat{v}_{j}=(\widehat{\upsilon}_{u_{1},(2),j},\dots,\widehat{\upsilon}_{u_{k},(2),j})^{\top}$. When $U^{\top}U$ is invertible (which holds with probability one for i.i.d. random directions as soon as $k\geq d$), the solution is
$\widehat{\mu}_{j}=(U^{\top}U)^{-1}U^{\top}\widehat{v}_{j}.$
Reconstruction of covariance matrices.
For each component $j=1,\dots,m$, we reconstruct $\Sigma_{j}\in\mathbb{S}^{+,d}$ by solving the convex quadratic program
$\widehat{\Sigma}_{j}:=\operatorname*{argmin}_{\Sigma\in\mathbb{S}^{+,d}}\ \sum_{r=1}^{k}\big(\langle u_{r},\Sigma u_{r}\rangle-\widehat{\sigma}^{2}_{u_{r},(2),j}\big)^{2}.$ (9)
As discussed in Proposition (Uniqueness of covariance reconstruction) in the main text, uniqueness holds if $\{u_{r}u_{r}^{\top}\}_{r=1}^{k}$ spans $\mathbb{S}^{d}$ (equivalently, if the associated design matrix has rank $d(d+1)/2$). In practice, (9) is solved with a semidefinite-programming routine (e.g. primal–dual interior-point methods, cf. [57, 81, 10]).
Output of Algorithm 1.
The final estimator is
$\widehat{\Theta}=\big(\widehat{\Lambda}_{(1)},\widehat{\bm{\mu}},\widehat{\bm{\Sigma}}\big),\qquad\widehat{\bm{\mu}}=(\widehat{\mu}_{1},\dots,\widehat{\mu}_{m}),\quad\widehat{\bm{\Sigma}}=(\widehat{\Sigma}_{1},\dots,\widehat{\Sigma}_{m}).$
Appendix B Example 3 (Section 3): Constrained-parameter estimation for a $d=20$ three-component Student $t$-mixture
We now consider a high-dimensional Student $t$-mixture in dimension $d=20$, and we compare the estimation obtained by the random projections approach (RP) with a multivariate EM procedure under a restricted parameterization. As in the previous examples, we generate i.i.d. observations from a finite mixture of multivariate $t$ distributions, but here we impose structural constraints on both the component locations and the scatter matrix in order to reduce the dimension of the parameter space.
Specifically, let $t_{\nu}(\mu,\Sigma)$ denote the $d$-variate Student $t$-distribution with $\nu$ degrees of freedom, location parameter $\mu\in\mathbb{R}^{d}$, and (scatter) matrix $\Sigma\in\mathbb{R}^{d\times d}$. We consider the three-component mixture
$F:=\lambda_{1}\,t_{\nu}(\mu_{1},\Sigma)+\lambda_{2}\,t_{\nu}(\mu_{2},\Sigma)+\lambda_{3}\,t_{\nu}(\mu_{3},\Sigma),\qquad\lambda_{1}+\lambda_{2}+\lambda_{3}=1,$
with common degrees of freedom $\nu=2$ and mixing weights
$(\lambda_{1},\lambda_{2},\lambda_{3})=(0.3,0.3,0.4).$
Constrained locations.
To reduce the location-parameter space, we constrain each mean vector to have constant coordinates,
$\mu_{k}=m_{k}\,\mathbf{1}_{d},\qquad k=1,2,3,$
where $\mathbf{1}_{d}$ denotes the $d$-dimensional vector of ones. In the simulation we set
$m_{1}=0,\qquad m_{2}=1,\qquad m_{3}=3,$
so that $\mu_{1}=(0,\ldots,0)$, $\mu_{2}=(1,\ldots,1)$, and $\mu_{3}=(3,\ldots,3)$.
Compound-symmetry covariance.
To reduce the scatter-parameter space, the three components share a common compound-symmetry (CS) matrix,
$\Sigma=\Sigma(x):=(1-x)I_{d}+x\,\mathbf{1}_{d}\mathbf{1}_{d}^{\top},$
that is, $\Sigma_{ii}=1$ and $\Sigma_{ij}=x$ for $i\neq j$. In the simulation we take $x=0.25$ (which must satisfy 1/(d1)<x<1-1/(d-1)<x<1 to ensure positive definiteness).
Sampling scheme.
For each replicate, we generate $N=200$ i.i.d. observations from $F$ by first sampling latent labels $Z_{i}\in\{1,2,3\}$ with $\mathbb{P}(Z_{i}=k)=\lambda_{k}$, and then drawing
$X_{i}\mid(Z_{i}=k)\sim t_{\nu}(\mu_{k},\Sigma),\qquad i=1,\ldots,n.$
This experiment is replicated $100$ times.
Estimation and evaluation.
In each replicate we estimate the mixture parameters under the above constraints using: (i) a multivariate EM algorithm that estimates $(\lambda_{1},\lambda_{2},\lambda_{3})$, $(m_{1},m_{2},m_{3})$ and $x$; and (ii) an RP-based procedure that combines (a) a screened collection of random projections, (b) one-dimensional $t$-mixture fits to recover $(m_{1},m_{2},m_{3})$ and $x$, and (c) a short multivariate refinement initialized at the RP estimates.
For both procedures, we compute MAP allocations and summarize clustering performance via the adjusted Rand index (ARI), see Figure 10. In addition, we measure estimation accuracy through absolute errors for the mixing weights, the mean scalars (reported as errors for $\mu_{1},\mu_{2},\mu_{3}$ under the constraint $\mu_{k}=m_{k}\mathbf{1}_{d}$), see Figures 11, 12, and 13; and the scatter matrix, see Figure 14.
Fig. 10: Boxplots of the Adjusted Rand Index (ARI) over the $100$ Monte Carlo replicates, comparing the EM and RP procedures. Larger ARI values indicate a better agreement between the estimated allocations and the true component labels.
Fig. 11: Boxplots of estimation errors over the $100$ Monte Carlo replicates for component $k=1$: (left) $|\mu_{1}-\mu_{1}^{*}|$ and (right) $|\lambda_{1}-\lambda_{1}^{*}|$, comparing EM and RP. Boxes show the median and interquartile range; whiskers follow Tukey’s rule (outliers omitted in the plot).
Fig. 12: Boxplots of estimation errors over the $100$ Monte Carlo replicates for component $k=2$: (left) $|\mu_{2}-\mu_{2}^{*}|$ and (right) $|\lambda_{2}-\lambda_{2}^{*}|$, comparing EM and RP. Boxes show the median and interquartile range; whiskers follow Tukey’s rule (outliers omitted in the plot).
Fig. 13: Boxplots of estimation errors over the $100$ Monte Carlo replicates for component $k=3$: (left) $|\mu_{3}-\mu_{3}^{*}|$ and (right) $|\lambda_{3}-\lambda_{3}^{*}|$, comparing EM and RP. Boxes show the median and interquartile range; whiskers follow Tukey’s rule (outliers omitted in the plot).
Fig. 14: Boxplots of the Frobenius error $\|\Sigma-\Sigma^{*}\|_{F}$ over the $100$ Monte Carlo replicates, comparing EM and RP.
Appendix C Descriptive Summary Tables and Confusion Matrices for the Simulations in Section 3
Table 1: Average values of the estimated parameters over $100$ replicates of the algorithm. Marginal standard deviations (sd) are shown in parentheses. Results are reported for four separability scenarios indexed by the location-shift parameter $\eta\in\{1/2,1,3/2,2\}$.
Algorithm Component $\mu_{1}$ $\mu_{2}$ $\sigma_{11}$ $\sigma_{22}$ $\sigma_{21}$
Average estimated values
Values 1 0 0 1 1/2 0
2 $\eta$ 0 1/2 1 0
Scenario: $\eta=1/2$
RP 1 -0.36 -0.04 0.92 0.90 0.00
sd (0.94) (1.30) (1.35) (0.65) (0.62)
2 0.92 0.04 0.66 1.02 -0.05
sd (0.65) (1.37) (0.59) (0.88) (0.48)
EM-st 1 0.03 -0.05 0.91 0.63 0.00
sd (1.04) (0.98) (2.61) (0.85) (0.90)
2 0.89 -0.01 0.50 0.64 0.01
sd (0.66) (0.69) (0.48) (0.83) (0.54)
Scenario: $\eta=1$
RP 1 -0.38 0.11 0.84 0.86 0.03
sd (0.71) (1.25) (0.48) (0.49) (0.33)
2 1.35 -0.01 0.57 0.80 0.02
sd (0.49) (1.07) (0.25) (0.45) (0.22)
EM-st 1 0.41 -0.01 0.55 0.50 -0.05
sd (0.91) (0.69) (0.49) (0.40) (0.50)
2 1.48 -0.01 0.71 0.79 -0.28
sd (0.98) (0.76) (2.65) (2.61) (2.65)
Scenario: $\eta=3/2$
RP 1 -0.14 -0.03 1.51 1.10 -0.36
sd (1.00) (1.12) (5.75) (3.47) (4.25)
2 1.73 -0.04 0.66 0.97 0.00
sd (0.45) (1.11) (0.31) (0.99) (0.45)
EM-st 1 0.87 0.13 1.12 0.75 0.17
sd (0.69) (0.97) (3.42) (0.90) (1.59)
2 1.39 0.10 0.44 0.64 -0.02
sd (0.65) (1.05) (0.28) (0.24) (0.30)
Scenario: $\eta=2$
RP 1 -0.27 0.01 1.10 0.65 -0.07
sd (0.39) (0.44) (0.43) (0.50) (0.34)
2 2.20 0.06 0.67 0.89 0.00
sd (0.27) (0.41) (0.28) (0.33) (0.17)
EM-st 1 0.89 -0.03 0.66 0.64 -0.09
sd (0.53) (0.88) (0.39) (0.33) (0.44)
2 1.60 -0.07 0.36 0.74 -0.02
sd (0.49) (1.02) (0.14) (0.26) (0.18)
$\eta$ EM-st RP
1/2 $\begin{pmatrix}0.163&0.139\\ 0.310&0.389\end{pmatrix}$ $\begin{pmatrix}0.133&0.169\\ 0.253&0.446\end{pmatrix}$
1.0 $\begin{pmatrix}0.081&0.188\\ 0.171&0.561\end{pmatrix}$ $\begin{pmatrix}0.150&0.118\\ 0.200&0.532\end{pmatrix}$
3/2 $\begin{pmatrix}0.203&0.093\\ 0.294&0.409\end{pmatrix}$ $\begin{pmatrix}0.198&0.099\\ 0.139&0.564\end{pmatrix}$
2.0 $\begin{pmatrix}0.156&0.141\\ 0.120&0.583\end{pmatrix}$ $\begin{pmatrix}0.206&0.091\\ 0.097&0.607\end{pmatrix}$
Table 2: Average confusion matrices (joint proportions) between true labels and posterior classifications for EM-st and RP.
Table 3: Average values of the estimated parameters over $100$ replicates of the algorithm. Marginal standard deviations (sd) are shown in parentheses. Results are reported for three contamination scenarios indexed by the outlier proportion $\gamma\in\{0.05,0.10,0.15\}$.
Algorithm Component $\mu_{1}$ $\mu_{2}$ $\sigma_{11}$ $\sigma_{22}$ $\sigma_{21}$
Average estimated values
Values 1 0 0 1 1/2 0
2 2 0 1/2 1 0
Scenario: $\gamma=0.05$
RP 1 -0.27 0.12 1.16 0.68 0.00
sd (0.32) (0.54) (0.41) (0.53) (0.31)
2 2.27 0.04 0.72 1.05 0.05
sd (0.40) (0.43) (0.45) (0.31) (0.12)
EM-rob 1 0.60 0.10 3.18 1.32 0.17
sd (0.32) (0.17) (0.75) (0.39) (0.43)
2 2.07 0.05 0.89 2.31 0.06
sd (0.10) (0.17) (0.37) (0.51) (0.16)
Scenario: $\gamma=0.10$
RP 1 -0.14 0.46 1.08 0.73 0.07
sd (0.45) (0.86) (0.27) (0.40) (0.20)
2 2.24 -0.11 0.68 1.09 0.05
sd (0.37) (0.68) (0.18) (0.40) (0.14)
EM-rob 1 0.70 0.17 3.33 1.61 0.34
sd (0.29) (0.23) (0.68) (0.66) (0.70)
2 2.07 0.13 0.93 2.53 0.06
sd (0.10) (0.23) (0.36) (0.74) (0.18)
Scenario: $\gamma=0.15$
RP 1 0.12 0.59 1.03 0.88 0.09
sd (0.63) (1.75) (0.35) (0.40) (0.24)
2 2.21 0.28 0.70 1.12 0.09
sd (0.42) (1.57) (0.25) (0.81) (0.21)
EM-rob 1 0.77 0.28 3.37 1.85 0.54
sd (0.31) (0.30) (0.73) (0.87) (0.84)
2 2.08 0.21 1.07 2.80 0.08
sd (0.11) (0.29) (0.36) (1.02) (0.22)
$\gamma$ RobEM RP
$\gamma=5\%$ $\begin{pmatrix}0.262&0.043\\ 0.157&0.538\end{pmatrix}$ $\begin{pmatrix}0.232&0.072\\ 0.063&0.632\end{pmatrix}$
$\gamma=10\%$ $\begin{pmatrix}0.254&0.046\\ 0.140&0.560\end{pmatrix}$ $\begin{pmatrix}0.191&0.109\\ 0.064&0.636\end{pmatrix}$
$\gamma=15\%$ $\begin{pmatrix}0.253&0.045\\ 0.178&0.524\end{pmatrix}$ $\begin{pmatrix}0.292&0.076\\ 0.103&0.529\end{pmatrix}$
Table 4: Average confusion matrices (joint proportions) between true labels and posterior classifications for RobEM-st and RP.